# Kang Liao1  Size Wu1  Zhonghua Wu2  Linyi Jin3

### v1

@@ Abstract
Camera-centric understanding and generation are two cornerstones of spatial intelligence, yet they are typically studied in isolation. We present Puffin, a unified camera-centric multimodal model that extends spatial awareness along the camera dimension. Puffin integrates language regression and diffusion-based generation to interpret and create scenes from arbitrary viewpoints. To bridge the modality gap between cameras and vision-language, we introduce a novel paradigm that treats camera as language, enabling thinking with camera. This guides the model to align spatially grounded visual cues with photographic terminology while reasoning across geometric context. Puffin is trained on Puffin-4M, a large-scale dataset of 4 million vision-language-camera triplets. We incorporate both global camera parameters and pixel-wise camera maps, yielding flexible and reliable spatial generation. Experiments demonstrate Puffin’s superior performance over specialized models for camera-centric generation and understanding. With instruction tuning, Puffin generalizes to diverse cross-view tasks such as spatial imagination, world exploration, and photography guidance. We will release the code, models, dataset pipeline, and benchmark to advance multimodal spatial intelligence research.
Figure 1: Illustration of the versatile capabilities of our Puffin model. It unifies camera-centric generation (a) and understanding (b), supports the thinking mode (c), and enables diverse applications (d).
@@ 1 Introduction
In this work, we make the first attempt to unify camera-centric understanding and generation in a cohesive framework. Motivated by recent progress in unified understanding and generation with large multimodal models (LMMs) (Team, 2024; Wu et al., 2025d; 2024; Pan et al., 2025; Wu et al., 2025b), we extend this paradigm to the spatial domain, where camera geometry plays a central role. However, unlike language or images, camera parameters are abstract and non-intuitive: they describe FoV, orientation, or perspective in numerical form rather than semantic content. This discrepancy introduces a modality gap when integrating cameras into LMMs. For instance, when users specify “20° roll” or “35mm lens” for controllable generation, existing models often ignore or misinterpret such cues, pursuing semantic alignment while neglecting precise spatial control. Similarly, current LMMs tend to collapse geometric details into coarse representations when understanding camera information, leading to spatially inconsistent outputs. As a result, naïvely extending LMMs cannot resolve conflicts between modalities, producing suboptimal performance in both tasks.
To address this challenge, we introduce Puffin, a unified multimodal framework that interprets cameras as a first-class modality. Puffin combines autoregressive and diffusion modeling to jointly perform camera-centric understanding and generation [^1] We mainly focus on single-view calibration and text-to-image controllable generation, but Puffin can be flexibly extended to cross-view understanding and generation via instruction tuning (see Figure 10).. Instead of treating camera parameters as auxiliary labels, Puffin introduces the notion of thinking with camera, aligning spatially grounded visual cues with professional photographic terminology while reasoning over geometric context. This design provides a shared chain-of-thought across multimodal tasks, enabling spatially consistent understanding and controllably aligned generation.
@@ 5.1 Implementation Details
Network Configuration. For the architecture of our Puffin, we use the pretrained C-RADIOv3-H (Heinrich et al., 2025), Qwen2.5-1.5B-Instruct (Qwen et al., 2024), and SD3-Medium (Esser et al., 2024) to initialize our geometry-aligned vision encoder, LLM, and diffusion model, respectively. Learnable queries with the number of 64 and a lightweight connector comprising six transformer layers are exploited to translate the LLM hidden states to conditioning signals for the diffusion model. The resolutions of the image and camera maps are set to $512\times 512$ for all tasks. For tokenization, the camera parameter tokenizer follows the same procedure as the text tokenizer. For the camera map, we adopt Perspective Field (Jin et al., 2023). We first normalize its values to the range [1,1][-1,1], and then reuse the image tokenizer (i.e., the VAE encoder), as the camera map also has three channels. Pretraining a specialized tokenizer for camera maps is left as future work.
@@
MegaDepth
DeepCalib (Lopez et al., 2019) 01.41 34.6 65.4 79.4 05.19 11.9 27.8 44.8 11.14 05.6 12.1 22.9
@@
TartanAir
DeepCalib (Lopez et al., 2019) 01.95 24.7 55.4 71.5 03.27 16.3 38.8 58.5 08.07 01.5 08.8 27.2
@@
LaMAR
DeepCalib (Lopez et al., 2019) 01.15 44.1 73.9 84.8 04.68 10.8 28.3 49.8 10.93 00.7 13.0 24.0
@@
Puffin-Und
DeepCalib (Lopez et al., 2019) 01.90 29.3 56.2 71.7 03.71 15.3 36.0 54.9 07.43 09.0 19.4 34.8
@@ 5.2 Evaluations on Camera Understanding
Approach Up Vector [degrees] Latitude [degrees] Gravity [degrees]
mean error $\downarrow$ median error $\downarrow$ mean error $\downarrow$ median error $\downarrow$ mean error $\downarrow$ median error $\downarrow$
GPT-4o* (OpenAI, 2025) 24.11 22.86 15.87 13.67 28.08 27.39
GPT-4o (OpenAI, 2025) 24.07 22.10 14.67 12.43 27.19 26.32
Qwen-Image* (Wu et al., 2025a) 23.80 22.73 15.76 13.90 27.75 27.22
Qwen-Image (Wu et al., 2025a) 23.98 22.60 15.92 13.92 27.86 26.45
Nano Banana* (Google DeepMind, 2025) 24.08 23.13 16.66 15.05 28.78 28.22
Nano Banana (Google DeepMind, 2025) 24.65 23.50 15.80 13.98 28.22 26.73
PreciseCam (Bernal-Berdun et al., 2025) 18.66 17.47 12.49 09.99 18.39 15.34
Puffin (Ours) 11.94 10.12 06.34 04.04 06.79 03.43
@@ 5.3 Evaluations on Camera-Controllable Generation
Settings. We evaluate our generation performance against the state-of-the-art method PreciseCam (Bernal-Berdun et al., 2025). In addition, we compare our approach with recent powerful multimodal models, including GPT-4o (OpenAI, 2025), Qwen-Image (Wu et al., 2025a), and Nano Banana (Google DeepMind, 2025), using the same captions as our method. The prompt templates are shown in Section A1.2. To mitigate the data gap for the multimodal models, we convert the camera parameters in captions from radians to degrees or express them using professional photographic terms. For quantitative evaluation, we adopt an offline method (Veicht et al., 2024) to estimate pixel-wise camera maps. Using the ground truth maps, we then compute the mean and median errors of the up vector, latitude, and gravity, all measured in degrees. Since no benchmark dataset exists for text-to-image generation with precise camera parameters, we construct Puffin-Gen to fill this gap. The dataset consists of 650 caption–camera pairs spanning diverse scenarios and camera configurations. We will release Puffin-Gen and Puffin-Und to support standardized evaluation and facilitate subsequent works.
@@ 5.5 Applications
We illustrate the versatile capabilities of our Puffin in Figure 10. Similar to previous methods (Hold-Geoffroy et al., 2018; Jin et al., 2023), Puffin can support virtual 3D object insertion into in-the-wild images by accurately predicting camera parameters. Furthermore, it can be flexibly extended to a range of cross-view tasks through instruction tuning, such as spatial imagination, world exploration, and photographic guidance. For both the initial and generated views in world exploration, we visualize 3D reconstruction results with VGGT (Wang et al., 2025a), showing proper spatial consistency across viewpoints. Additional results are presented in Figure A7, A8, A9.
@@ Acknowledgment
This research is supported by cash and in-kind funding from NTU S-Lab and industry partner(s). We thank Zhouxia Wang, Zongsheng Yue, Haiwen Diao, Zhaoxi Chen, Zhijie Shen, Xiang Li, Qiqi Gong, Edurne Bernal-Berdun, and Koki Maeda for their insightful discussions. We thank Xu Song and Juncheng Zhou for their help with the comparison methods.
@@
Term ($t$)
Large counterclockwise
Dutch angle
Small counterclockwise
Dutch angle
Near level shot
Small clockwise
Dutch angle
Large clockwise
Dutch angle
@@
Stanford2D3D
DeepCalib (Lopez et al., 2019) 01.59 33.8 63.9 79.2 02.58 21.6 46.9 65.7 06.67 08.1 20.6 37.6
@@
A2.3 Downstream Applications

### v2

@@ Abstract
Camera-centric understanding and generation are two cornerstones of spatial intelligence, yet they are typically studied in isolation. We present Puffin, a unified camera-centric multimodal model that extends spatial awareness along the camera dimension. Puffin integrates language regression and diffusion-based generation to interpret and create scenes from arbitrary viewpoints. To bridge the modality gap between cameras and vision-language, we introduce a novel paradigm that treats camera as language, enabling thinking with camera. This guides the model to align spatially grounded visual cues with photographic terminology while reasoning across geometric context. Puffin is trained on Puffin-4M, a large-scale dataset of 4 million vision-language-camera triplets. We incorporate both global camera parameters and pixel-wise camera maps, yielding flexible and reliable spatial generation. Experiments demonstrate Puffin’s superior performance over specialized models for camera-centric generation and understanding. With our designed instruction tuning, Puffin generalizes to diverse cross-view tasks such as spatial imagination, world exploration, and photography guidance. To advance multimodal spatial intelligence research, we released the code, models, dataset pipeline, and benchmark at https://kangliao929.github.io/projects/puffin.
Figure 1: Illustration of the versatile capabilities of our Puffin model. It unifies camera-centric generation (a) and understanding (b), supports the thinking mode (c), and enables diverse 3D applications (d).
@@ 1 Introduction
In this work, we make the first attempt to unify camera-centric understanding and generation in a cohesive framework. Motivated by recent progress in unified understanding and generation with large multimodal models (LMMs) (Team, 2024; Wu et al., 2025d; 2024; Pan et al., 2025; Wu et al., 2025b), we extend this paradigm to the spatial domain, where camera geometry plays a central role. However, unlike language or images, camera parameters are abstract and non-intuitive: they describe field-of-view (FoV), orientation, or perspective in numerical form rather than semantic content. This discrepancy introduces a modality gap when integrating cameras into LMMs. For instance, when users specify “20° roll” or “35mm lens” for controllable generation, existing models often ignore or misinterpret such cues, pursuing semantic alignment while neglecting precise spatial control. Similarly, current LMMs tend to collapse geometric details into coarse representations when understanding camera information, leading to spatially inconsistent outputs. As a result, naïvely extending LMMs cannot resolve conflicts between modalities, producing suboptimal performance in both tasks.
To address this challenge, we introduce Puffin, a unified multimodal framework that interprets cameras as a first-class modality. Puffin combines autoregressive and diffusion modeling to jointly perform camera-centric understanding and generation [^1] We mainly focus on single-view calibration and text-to-image controllable generation, but Puffin can be flexibly extended to cross-view understanding and generation via our designed instruction tuning (see Figure 10).. Instead of treating camera parameters as auxiliary labels, Puffin introduces the notion of thinking with camera, aligning spatially grounded visual cues with professional photographic terminology while reasoning over geometric context. This design provides a shared chain-of-thought across multimodal tasks, enabling spatially consistent understanding and controllably aligned generation.
@@ 5.1 Implementation Details
Network Configuration. For the architecture of our Puffin, we use the pretrained C-RADIOv3-H (Heinrich et al., 2025), Qwen2.5-1.5B-Instruct (Qwen et al., 2024), and SD3-Medium (Esser et al., 2024) to initialize our geometry-aligned vision encoder, LLM, and diffusion model, respectively. Learnable queries with the number of 64 and a lightweight connector comprising six transformer layers are exploited to translate the LLM hidden states to conditioning signals for the diffusion model. The resolutions of the image and camera maps are set to $512\times 512$ for all tasks. For tokenization, the camera parameter tokenizer follows the same procedure as the text tokenizer. Since camera parameters are numerical, we first serialize them into discrete tokens, which are naturally handled by the standard text tokenizer without requiring any extra module. Introducing an additional tokenizer (or a separate text encoder) would substantially increase the alignment burden across modules and modalities in a unified multimodal model. For this reason, we keep the vanilla decoder-only LLM backbone and its tokenizer to process both language and camera parameters. For the camera map, we adopt Perspective Field (Jin et al., 2023). We first normalize its values to the range [1,1][-1,1], and then reuse the image tokenizer (i.e., the VAE encoder), as the camera map also has three channels. Pretraining a specialized tokenizer for camera maps is left as future work.
@@
MegaDepth DeepCalib (Lopez et al., 2019) 01.41 34.6 65.4 79.4 05.19 11.9 27.8 44.8 11.14 05.6 12.1 22.9
@@
TartanAir DeepCalib (Lopez et al., 2019) 01.95 24.7 55.4 71.5 03.27 16.3 38.8 58.5 08.07 01.5 08.8 27.2
@@
LaMAR DeepCalib (Lopez et al., 2019) 01.15 44.1 73.9 84.8 04.68 10.8 28.3 49.8 10.93 00.7 13.0 24.0
@@
Puffin-Und DeepCalib (Lopez et al., 2019) 01.90 29.3 56.2 71.7 03.71 15.3 36.0 54.9 07.43 09.0 19.4 34.8
@@ 5.2 Evaluations on Camera Understanding
Approach Up Vector [degrees] Latitude [degrees] Gravity [degrees] Visual Quality
mean error $\downarrow$ median error $\downarrow$ mean error $\downarrow$ median error $\downarrow$ mean error $\downarrow$ median error $\downarrow$ FID $\downarrow$
GPT-4o* (OpenAI, 2025) 24.11 22.86 15.87 13.67 28.08 27.39 95.92
GPT-4o (OpenAI, 2025) 24.07 22.10 14.67 12.43 27.19 26.32 94.43
Qwen-Image* (Wu et al., 2025a) 23.80 22.73 15.76 13.90 27.75 27.22 83.31
Qwen-Image (Wu et al., 2025a) 23.98 22.60 15.92 13.92 27.86 26.45 83.37
Nano Banana* (Google DeepMind, 2025) 24.08 23.13 16.66 15.05 28.78 28.22 91.66
Nano Banana (Google DeepMind, 2025) 24.65 23.50 15.80 13.98 28.22 26.73 88.02
PreciseCam (Bernal-Berdun et al., 2025) 18.66 17.47 12.49 09.99 18.39 15.34 90.91
Puffin (Ours) 11.94 10.12 06.34 04.04 06.79 03.43 69.46
@@ 5.3 Evaluations on Camera-Controllable Generation
Settings. We evaluate our generation performance against the state-of-the-art method PreciseCam (Bernal-Berdun et al., 2025). In addition, we compare our approach with recent powerful multimodal models, including GPT-4o (OpenAI, 2025), Qwen-Image (Wu et al., 2025a), and Nano Banana (Google DeepMind, 2025), using the same captions as our method. The prompt templates are shown in Section A1.2. To mitigate the data gap for the multimodal models, we convert the camera parameters in captions from radians to degrees or express them using professional photographic terms. For quantitative evaluation, we adopt an offline method (Veicht et al., 2024) to estimate pixel-wise camera maps. Using the ground truth maps, we then compute the mean and median errors of the up vector, latitude, and gravity, all measured in degrees. We also incorporate the standard FID metric to assess the overall visual quality of the generated images. Since no benchmark dataset exists for text-to-image generation with precise camera parameters, we construct Puffin-Gen to fill this gap. The dataset consists of 650 caption–camera pairs spanning diverse scenarios and camera configurations. We will release Puffin-Gen and Puffin-Und to support standardized evaluation and facilitate subsequent works.
@@ 5.4 Ablation Studies
For other general-purpose unified models, the understanding tasks mainly target high-level concepts such as recognition and semantic comprehension. As a result, the domain gaps across multimodal tasks are more pronounced in these models, likely requiring more delicate architectural designs to harmonize representations across different modalities.
@@ 5.5 Applications
We illustrate the versatile capabilities of our Puffin in Figure 10. Similar to previous methods (Hold-Geoffroy et al., 2018; Jin et al., 2023), Puffin can support virtual 3D object insertion into in-the-wild images by accurately predicting camera parameters. Furthermore, it can be flexibly extended to a range of cross-view tasks through instruction tuning, such as spatial imagination, world exploration, and photographic guidance. For both the initial and generated views in world exploration, we visualize 3D reconstruction results with VGGT (Wang et al., 2025a), showing proper spatial consistency across viewpoints. Additional results are presented in Figure A9, A10, A11.
@@ Acknowledgment
This study is supported under the RIE2020 Industry Alignment Fund Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). It is also supported by Singapore MOE AcRF Tier 2 (MOE-T2EP20224-0003). We thank Zhouxia Wang, Zongsheng Yue, Haiwen Diao, Zhaoxi Chen, Zhijie Shen, Xiang Li, Qiqi Gong, Edurne Bernal-Berdun, and Koki Maeda for their insightful discussions. We thank Xu Song and Juncheng Zhou for their help with the comparison methods.
@@
Term ($t$) Large counterclockwise Dutch angle Small counterclockwise Dutch angle Near level shot Small clockwise Dutch angle Large clockwise Dutch angle
@@
Stanford2D3D DeepCalib (Lopez et al., 2019) 01.59 33.8 63.9 79.2 02.58 21.6 46.9 65.7 06.67 08.1 20.6 37.6
@@
A2.3 Analysis
Table A3: Model size comparisons: the specialized understanding and generation models (GeoCalib (Veicht et al., 2024) and PreciseCam (Bernal-Berdun et al., 2025)), and the proposed unified camera-centric model.
Model Type Parameters GFLOPs
GeoCalib (Veicht et al., 2024) Specialized Model (Understanding) 28.9M $3.47\times 10^{2}$
PreciseCam (Bernal-Berdun et al., 2025) Specialized Model (Generation) 01.3B $2.67\times 10^{3}$
Puffin-4M (Ours) Unified Model 04.4B $2.92\times 10^{5}$
Model Size. We show the comparison results on the total parameter count and FLOPs with previous understanding and generation models in Table A3, such as GeoCalib Veicht et al. (2024) and PreciseCam Jin et al. (2023). While Puffin is larger than previous specialized models, it replaces separate understanding and generation networks with a single unified model that handles both tasks within one framework. This design not only simplifies deployment, but also allows us to fully exploit a high-capacity backbone when training on large-scale multimodal datasets. In terms of overall parameter count and FLOPs, Puffin remains substantially more affordable than recent general-purpose unified multimodal models such as Bagel (14B) Deng et al. (2025) and Qwen-Image (20B) Wu et al. (2025a).
Error Analysis. In addition to the metrics (error and AUC) used in the main experiments, we also add the error bars for the representative baselines and our method in Figure A2(a) to provide a clearer and more comprehensive comparison. The results show that Puffin exhibits better robustness across the entire data distribution. The improvement is consistent across all intrinsics/extrinsics components and remains stable even under challenging camera configurations.
Model vs. Data. We conduct experiments where the comparison method (Veicht et al., 2024) is re-trained on the same dataset (Puffin-4M) as ours, strictly following its original training recipe. Interestingly, we find that the re-trained GeoCalib (Veicht et al., 2024) on our 4M dataset slightly underperforms the original GeoCalib model trained on its 40K dataset. The detailed evaluation results on the Puffin-Und test set are reported in Table A4.
Table A4: Additional evaluation results on camera understanding by re-training the comparison method (Veicht et al., 2024) on our constructed Puffin-4M dataset*.
Approach Roll [degrees] Pitch [degrees] FoV [degrees]
GeoCalib* (Veicht et al., 2024) 01.12 46.0 69.3 80.1 02.54 24.5 47.5 65.4 05.47 10.8 25.5 43.8
[…]
Figure A2: More data analysis of the experiments. (a) We show the error bar of the camera understanding evaluation on GeoCalib (Veicht et al., 2024) and Puffin. (b) We show how the performance of Puffin scales with data size across different camera parameters.
In contrast, Puffin is built on a high-capacity LLM backbone, which, like other large multimodal models (e.g., LLaVA, Qwen-VL, InternVL), requires sufficiently large and diverse training data to avoid overfitting and to learn an accurate joint distribution over images, camera geometry, and language. In this sense, Puffin-4M is not merely a “bonus”, but rather a necessary data regime for such a unified multimodal model that jointly supports both understanding and generation.
Furthermore, we also conduct additional experiments by re-training our model on Puffin-4M of different scales (Stage-II SFT). As shown in Figure A2 (b), we observe clear and consistent improvements as the data scale increases. Overall, these results suggest that the dataset and the model contribute jointly and should not be viewed in isolation.
Beyond the above scaling results in Figure A2 (b), we find an interesting difference in trends across camera parameters. For roll, the model learns quickly even with relatively limited data, since it mainly relies on low- to mid-level geometric cues that are easy to capture (e.g., strongly slanted lines indicating a large roll). By contrast, estimating pitch and FoV requires more holistic and high-level spatial understanding, which cannot be sufficiently captured by local visual patterns alone and therefore benefits more from larger-scale data to form robust spatial reasoning concepts. This observation is consistent with our discussion in Section A2.1. Based on these trends, we believe that further scaling the dataset would bring additional gains in camera geometry understanding, especially for pitch and FoV.
Multi-turn Interleaved Capability. Exploring the multi-round conversational capability of a unified multimodal model is meaningful. To this end, we conduct experiments on multi-turn interleaved conversations (generation $\rightarrow$ understanding and understanding $\rightarrow$ generation). As shown in Figure A3, Puffin can carry out coherent interleaved dialogues conditioned on previous turns. Specifically, it produces consistent cross-view generation results based on its previous reasoning, and accurate camera understanding based on its own generated images. This demonstrates that Puffin not only supports both capabilities within a single framework but can also use them in an interactive way over multiple conversational rounds, without any task-specific switching or separate models.
Figure A3: Multi-turn interleaved capability of Puffin.
A2.4 Downstream Applications